# Strong and noise-tolerant entanglement in dissipative optomechanics

### v2

@@ V.2 Purely dissipative coupling ($g_{\omega}=0$)
V.3 Combined coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
To examine the interplay between dispersive and dissipative coupling effects on optomechanical entanglement, we set the dispersive coupling strength to $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the purely dispersive coupling regime [refer to Fig. 3(a)]. By progressively increasing the dissipative coupling strength, we observe a gradual decline in the entanglement, as shown in Fig. 5(a). The reduction in entanglement possibly arises from the interference between the dispersive and dissipative couplings. When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig. 5(b). Moreover, the larger $g_{\kappa}$ is not benificial to entanglement generation. This can also be revealed by Fig. 5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are taken . For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K. But for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can be up to $8$ K.
@@ VI Conclusion
In this paper, we investigate quantum entanglement in a dissipative optomechanical system, acted by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the BS, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case. Our analysis reveals that, in the purely dispersive case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the purely dissipative case, the displacement shows a linear dependence. Furthermore, entanglement generated through purely dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via purely dispersive coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest the potential of purely dissipative coupling to engineer strong and noise-tolerant quantum entanglement, offering a promising pathway for advancing quantum technologies.

### v3

@@ V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
To examine the effect of the cooperative coupling (i.e., coherent and dissipative couplings) on optomechanical entanglement, we set the coherent coupling strength to be $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the dispersive coupling case [refer to Fig. 3(a)]. By progressively increasing the dissipative coupling strength, we observe a gradual decline in the entanglement, as shown in Fig. 5(a). The reduction in entanglement possibly arises from the interference between the coherent and dissipative couplings. When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig. 5(b). Moreover, the larger $g_{\kappa}$ is not benificial to entanglement generation. This can also be revealed by Fig. 5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are taken . For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K. But for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can be up to $8$ K.
@@ VI Conclusion
In a summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling case, dissipative case, and cooperative case. Our results reveals that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the dissipative coupling case, the displacement shows a linear dependence. Furthermore, entanglement generated through dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via coherent coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest that dissipative coupling may be a promising mechanism to engineer strong and noise-tolerant quantum entanglement.